OpenStudy (stbest):

Will Give MEDAL im writing a quiz correction but idk how i got this wrong What are the coordinates of the center of the circle that passes through the points (1,1)(1,5)(5,5) (0,0) (2.5,2.5) Correct Answer (3,3) Incorrect Response (5,1) PLZ EXPLAIN

OpenStudy (stbest):

no pic in this one

OpenStudy (stbest):

i thought it was 5,1 because they were the only ones that were in the question i didn't even think of 3,3

OpenStudy (stbest):

1 More

OpenStudy (stbest):

explain that 2

OpenStudy (anonymous):

let the eq. of circle be where centre is (-g,-f) \[x^2+y^2+2gx+2fy+c=0 ...(1)\] because (1,1),(1,5),(5,5) lie on it. 1+1+2g+2f+c=0 or 2g+2f+c=-2 ... (2) 1+25+2g+10f+c=0 or 2g +10f+c =-26 ...(3) 25 +25+10 g+10 f+c=0 10 g+10 f+c=-50 ...(4) subtract (3) from (2) -8f=24 g=-3 subtract (4) from (2) -8g-8f=48 g+f=-6 -3+f=-6 f=-6+3=-3 centre is (3,3)

OpenStudy (stbest):

WOW THANKS !!!!